A rope so lies on the table that part of it lays over. the rope begins to slide when the length of the hanging part is 25% of entire length. the coefficient of friction between the rope and table is

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
k=1/3

l/4*mg=k*3l/4 mg